OpenStudy (anonymous):

Can someone please tell me what I am doing wrong? the question is Solve the following initial-value problem 1y''+1y=sin(x);y(0)=9y(0)=−6 y=

OpenStudy (anonymous):

My work r^2+1=0-->r==i,-i y_c=C1cosx+c2sinx f(x)=sinx y_p(x)=Asinx y'p(x)=Acosx y''p(x)=-sinx y''+y=sinx -Asinx+Asinx=sinx

OpenStudy (anonymous):

That cannot be true because I'm not supposed to get zero for the left hand side

OpenStudy (anonymous):

oh and btw its supposed to read y(0)=9 and y'(0)=-6

OpenStudy (anonymous):

You suppose: yp=Asinx then yp''=Asinx so: yp''+yp=2Asinx Must be equal to: sinx so: 2Asinx=sinx 2A=1 A=1/2 So: yp=(1/2)*sinx (this is the particular solution) But now we have to find the complementary one. we set: yh=c1*cos(x)+c2*sin(x) So now the solution that we are looking for is like this: y=c1*cosx+c2*sinx+(1/2)*sin(x) And now we have to use the initial conditions.Ok? Do you remember how do continue?

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

wait, im sorry but I'm confused on what you did! y''=-Asinx so that would be -Asinx+Asinx=sinx but because there's two of the same it becomes -Asinx+Asinx=2sinx but theres a negative, does that mean we get rid of the negative

OpenStudy (anonymous):

but then when you plug the values in, it doesnt make sense because I get yp=(1/2)sinx y'=(1/2)cosx, if I were to get rid of the negative

OpenStudy (anonymous):

but then when you put y(0)=9 and y'(0)=-6, you get an answer that doesnt make any sense

OpenStudy (unklerhaukus):

\[y_h=c_1\cos(x)+c_2\sin(x)\] \[y_p≠A\sin(x)\]

OpenStudy (anonymous):

I'm sorry but I think I am misunderstanding what it is were supposed to be doing

OpenStudy (anonymous):

so we derive y_h?

OpenStudy (unklerhaukus):

if the homogenous solution already has the form of the 'guess' of the y_p form in undetermined coefficients , there is resonance so multiply by x \[y_p=x\left(A\sin (x)+B\cos (x)\right)\]

OpenStudy (anonymous):

wait i think I got it! the general solution is yx=yh+yp and e found the particular solution to be (1/2)sin x so plus that in and hat we found to be the complimentary solution and we derive the general solution and then find the values of c1 and c2 using the values given. right?

OpenStudy (anonymous):

so youd getc1=9 and c2=2.5

OpenStudy (anonymous):

so the solution would be y(x)=9cosx+2.5sinx+.5sinx

OpenStudy (anonymous):

Be careful, the particular is not Asin(x), it's (-1/2)*x*cos(x) You have to use what UnkleRhaukus said After you have the particular, you add it to the complementary. And then use the initial conditions. Makes sense?

OpenStudy (anonymous):

so yp=x(asinx+Bcosx

OpenStudy (anonymous):

but why?

OpenStudy (anonymous):

how does that ebcome (-.5)xcosx?

OpenStudy (anonymous):

become*

OpenStudy (amistre64):

you can use wroniskian, or use the yc parts in a variation of parameters is my thought

OpenStudy (anonymous):

Thats what I would use to but the directions say otherwise! Thanks though!

OpenStudy (anonymous):

I understand everything but the particular solution part

OpenStudy (amistre64):

\[y_c=c_1cosx+c_2sinx\] \[y_p=A(x)cosx+B(x)sinx\] \[y'_p=A'(x)cosx+B'(x)sinx\]\[\hspace{3em}-A(x)sinx+B(x)cosx\] the A'B' parts = 0 to root out the homogenous part \[y'_p=-A(x)sinx+B(x)cosx\] \[ y''_p=-A'(x)sinx+B'(x)cosx\]\[\hspace{3em} -A(x)cosx-B(x)sinx\] y''p+ yp = sinx \[\ \ \ y_p=\cancel{A(x)cosx+B(x)sinx}\]\[ y''_p=-A'(x)sinx+B'(x)cosx\]\[\hspace{3em} \cancel{-A(x)cosx-B(x)sinx}\]--------------------------- \[ \hspace{3em} -A'(x)cosx +B'(x)sinx=sinx\] we have a system of 2 equations in 2 unknowns: A' and B' \[-A'(x)sinx+B'(x)cosx=0\]\[\ \ A'(x)cosx +B'(x)sinx=sinx\] determine the values of A' and B' then integrate them back into A and B the Wronskian just goes straight to the last part

OpenStudy (anonymous):

ok thanks sooo much!!

OpenStudy (amistre64):

youre welcome, i find the longer method to be more informative since the shorter methods suppose that you already have a working understanding